OpenStudy (anonymous):

A group of college students are volunteering for Help the Homeless during their spring break. They are putting the finishing touches on a house they built. Working alone, Jillian can paint a certain room in 14 hours. Mike can paint the same room in 3 hours. Write an equation that can be used to find how long it will take them working together to paint the room.

OpenStudy (anonymous):

the thing to remember in doing these problems is 1) the whole job is 1. 1 job 2) the rate is 1 over number of hours needed. so the equation here is \[(\frac{1}{14}+\frac{1}{3})x=1\] or if you prefer \[\frac{17}{42}x=1\] or even \[17x=42\]

OpenStudy (anonymous):

Still need help here?